# Reflexive Guidance: Improving OoDD in Vision-Language Models via Self-Guided Image-Adaptive Concept Generation

### v1

@@ Abstract
With the recent emergence of foundation models trained on internet-scale data and demonstrating remarkable generalization capabilities, such foundation models have become more widely adopted, leading to an expanding range of application domains. Despite this rapid proliferation, the trustworthiness of foundation models remains underexplored. Specifically, the out-of-distribution detection (OoDD) capabilities of large vision-language models (LVLMs), such as GPT-4o, which are trained on massive multi-modal data, have not been sufficiently addressed. The disparity between their demonstrated potential and practical reliability raises concerns regarding the safe and trustworthy deployment of foundation models. To address this gap, we evaluate and analyze the OoDD capabilities of various proprietary and open-source LVLMs. Our investigation contributes to a better understanding of how these foundation models represent confidence scores through their generated natural language responses. Based on our observations, we propose a self-guided prompting approach, termed Reflexive Guidance (ReGuide), aimed at enhancing the OoDD capability of LVLMs by leveraging self-generated image-adaptive concept suggestions. Experimental results demonstrate that our ReGuide enhances the performance of current LVLMs in both image classification and OoDD tasks.
@@ 1 Introduction
Thanks to substantial advancements in hardware and the availability of large-scale datasets far exceeding conventional ones, foundation models have achieved remarkable performance across a wide range of tasks, demonstrating exceptional generalization. This evolution has shifted toward leveraging multiple data modalities, particularly vision and natural language. As a result, vision-language foundation models have demonstrated their capabilities across diverse domains, from general tasks like text-to-image generation and vision-question answering, to specialized fields such as medical diagnosis and robotics (Esser et al., 2024; Li et al., 2023; Wake et al., 2023; Majumdar et al., 2024).
Despite the widespread adoption of these large vision-language models (LVLMs) including GPT, Claude, and Gemini (OpenAI, 2024; Anthropic, 2024; Reid et al., 2024), their trustworthiness and reliability have not been adequately investigated. Ensuring the robustness of deep neural networks has been a key research area to guarantee their safe application in practice. With the rapid popularization of LVLMs, practical concerns such as harmful content filtering and domain generalization have been actively studied (Zhang et al., 2024b; Han et al., 2024). However, fundamental aspects, such as the quality of confidence estimates for model predictions—extensively studied in single-modal models—remain underexplored. Some works have studied OoDD in CLIP (Ming et al., 2022; Jiang et al., 2024; Cao et al., 2024) and Stable Diffusion (Zhu et al., 2024), but few have examined OoDD in LVLMs, particularly with respect to confidence scores expressed through their natural language responses. This gap between their demonstrated capabilities and real-world reliability raises concerns about ensuring the safe and dependable deployment of LVLMs.
To bridge the gap, we first evaluate and compare the OoDD capabilities of LVLMs. Due to the lack of prior experimental configurations, we develop a framework for evaluating the OoDD capabilities of LVLMs. Our focus is on detecting OoD image inputs based on the in-distribution (ID) space specified by text inputs. To this end, we design a prompt that defines the ID space and guides the LVLM to provide confidence estimates through its generated responses. Using this prompt, we evaluate the OoDD capabilities of both proprietary and open-source LVLMs and analyze their behavior in expressing confidence from various perspectives. Fig. 1(a) summarizes the comparison of the evaluated LVLMs. In general, proprietary models outperform open-source models, and open-source models show performance improvements as model size increases. However, some open-source models including LLaVa-v1.6 (Mistral-7B) (Li et al., 2024a) and GLM-4v-9B (GLM et al., 2024) exhibit low OoDD performance despite achieving decent results on popular VLM benchmarks. [^1] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard We found that one of the reasons for this discrepancy is the insufficient image interpretation capabilities of these models, which limits their ability to accurately distinguish between ID and OoD inputs.
To enhance the OoD detectability of LVLMs, we propose a two-stage self-guided prompting approach called Reflexive Guidance (ReGuide). In the first stage, ReGuide prompts an LVLM to suggest two groups of concepts based on the given image: semantically similar concepts (i.e., near-OoD) and semantically dissimilar concepts (i.e., far-OoD). In the second stage, these suggested concepts are employed as auxiliary OoD classes. To the best of our knowledge, this is the first study to leverage image inputs to generate informative texts for OoDD. By utilizing the visual interpretation capabilities of LVLMs, ReGuide remains simple and model-agnostic, allowing the same prompt to be applied to different LVLMs. Fig. 1(b) shows the effectiveness of ReGuide. Notably, ReGuide significantly boosts the overall performance of open-source models, making them comparable to proprietary models. GPT-4o, which is the top performer prior to applying ReGuide, also benefits from ReGuide, especially on near-OoD datasets (e.g., NINCO, SSB Hard). Our results highlight the effectiveness of guiding LVLMs through self-generated, image-adaptive concept suggestions for OoDD.
@@
Figure 1: Overall performance of proprietary and open-source LVLMs on the ImageNet200 benchmark: (a) state-of-the-art LVLMs comparison and (b) the boosting effect of our ReGuide
@@ 2 Related work
In single-modal vision models, OoDD has been actively studied. Starting with the baseline method of using the maximum softmax value as an OoD score (Hendrycks & Gimpel, 2017), methods have evolved to improve confidence modeling (Moon et al., 2020; Liu et al., 2020; Bibas et al., 2021), post-hoc techniques (Liang et al., 2018; Lee et al., 2018b; Sun et al., 2021; Djurisic et al., 2023; Sun et al., 2022; Zhang et al., 2023; Liu & Qin, 2024), or a combination of both (Xu et al., 2024). Another approach involves leveraging auxiliary OoD samples to better distinguish between ID and OoD inputs. While real OoD samples show strong detection performance (Hendrycks et al., 2019; Chen et al., 2021; Lu et al., 2023; Katz-Samuels et al., 2022; Koo et al., 2024), they require access to real OoD data. Synthetic samples offer an alternative, providing the benefits of auxiliary OoD data without the need for collecting real samples (Lee et al., 2018a; Du et al., 2022; Tao et al., 2023; Zheng et al., 2023).
@@ 3.1 Problem Definition
We frame the OoDD problem for LVLMs based on the zero-shot OoDD scenario defined for CLIP. Similarly, $\mathcal{Y}$ is set to a fixed word set containing only ID class words. Given both $\mathbf{x}$ and $\mathcal{Y}$ as inputs, an LVLM is instructed to produce prediction results in a structured format. Then, class predictions and confidence estimates for $\mathbf{x}$ are extracted from the natural language responses by the LVLM. If the ground-truth label of $\mathbf{x}$ is not in $\mathcal{Y}$, the LVLM should provide low confidence scores for all $y_{i}$ in $\mathcal{Y}$ in its answers.
@@
3.2 Prompt Design
Fig. 2 illustrates the framework for OoDD evaluation on LVLMs with a simplified prompt. Our prompt consists of four components: a task description, an explanation of the rejection class, guidelines, and examples for the response format. Unlike previous work on LVLMs (Han et al., 2024; Hwang et al., 2024; Groot & Valdenegro-Toro, 2024), we encountered a significant number of failure cases when using a simple prompt consisting of the task statement including $\mathcal{Y}$ and a formatted output structure. We attribute this inconsistency to differences in how confidence scores are assigned in our framework. In our prompt design, the LVLM is expected to provide confidence scores for each class in $\mathcal{Y}$, whereas prior works require only a single confidence score for the predicted class. To mitigate these failures, we enhance the prompt by adding the following components, resulting in the final prompt used in our experiments. The complete prompt can be found in Appendix B.6.
Rejection class. When we provide only $\mathcal{Y}$, predictions for OoD samples often do not correspond to any class in $\mathcal{Y}$. Due to its strong zero-shot visual recognition capabilities, the LVLM can either provide the ground-truth label for an OoD input image or reject classification into any of the classes in $\mathcal{Y}$. To address this unintended behavior, we introduce a rejection class named none of these classes and provide a detailed explanation of its role to the LVLM.
Guidelines addressing failure cases. In addition to the case where the prediction does not fall within $\mathcal{Y}$, we identified several other issues, such as mismatches between the class assigned the highest confidence score and the predicted class, assigning a confidence score of 0.0 to all classes including the rejection class, or failing to provide confidence scores for all classes in $\mathcal{Y}$. To reduce the frequency of these failures, we include guidelines that address the most common cases.
Response examples. In-context learning (ICL) is commonly used to improve response quality by providing models with examples that illustrate the desired format during inference, allowing them to generate outputs that better align with these examples (Brown et al., 2020). We leverage ICL to reduce failure caused by formatting issues. However, we observed that when only a single example is provided, the LVLM tends to mimic the example, regardless of whether the input image is ID or OoD. To mitigate this tendency, we provide examples for both ID and OoD input images.
@@ 3.4 Experimental Settings
Comparison models. To compare LVLMs from diverse perspectives, we consider both proprietary and open-source models. For proprietary models, we employ three state-of-the-art (SOTA) models: GPT-4o (2024-08-06) (OpenAI, 2024), Gemini Pro 1.5 (Reid et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024). For open-source models, we use five models: LLaVA-v1.6-Mistral-7B (LLaVA-v1.6) (Li et al., 2024a), GLM-4v-9B (GLM-4v) (GLM et al., 2024), QWEN-VL-Chat (QWEN) (Bai et al., 2023), InternVL2-InternLM2-Chat-26B (InternVL2-26B), and InternVL2-LLaMA3-76B (InternVL2-76B) (Chen et al., 2024). Additionally, we include OpenCLIP (ViT-B-32 pretrained on LAION 2B-s34b-b79k) (Cherti et al., 2023) as a non-generative LVLM. To further facilitate comparison between single- and multi-modal models, we also include three single-modal state-of-the-art vision OoDD models, SCALE (Xu et al., 2024), fDBD (Liu & Qin, 2024), AugMix+ASH (Djurisic et al., 2023; Hendrycks et al., 2020).
@@
Evaluation metrics. We measure the OoDD performance using two commonly used metrics in OoDD: the area under the receiver operating characteristic curve (AUROC) and the false positive rate at the 95% true positive rate (FPR@95%TPR; FPR). We also include the ratio and number of valid responses—those that adhere to the provided guidelines and example format—as part of the evaluation metrics. We consider the ability to understand given instructions to be one of the model’s key capabilities. Only valid responses are included when measuring performance.
Table 1: Comparison results on the ImageNet200 benchmark. Full model names are provided in the footnotes. ‘Valid’ indicates the ratio of valid responses out of a total of $23,031$ image-prompt queries, with the number in brackets showing the valid response count. Bold highlights the best performance among generative LVLMs.
@@ 3.5 Results
Tab. 1 presents the OoDD capabilities of the compared models on the ImageNet200 benchmark. The near- and far-OoD results for SCALE and fDBD represent the average performance across their respective categories. ‘All OoD’ refers to the performance in distinguishing all OoD inputs, including near- and far-OoD inputs, from ID inputs (i.e., ID vs. all OoD). The results of QWEN-VL-Chat are omitted due to its exceptionally low ability to follow instructions, with a valid response rate of less than $1\%$.
@@
On both image recognition and OoDD tasks, proprietary models outperform open-source models in most cases, with reasonable valid response rates. All compared models have more difficulty in detecting near-OoD than far-OoD. The overall performance ranking of the evaluated LVLMs generally aligns with the OpenVLM leaderboard [^3] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard, except for Gemini Pro 1.5. In our results, Gemini Pro 1.5 shows better ID accuracy and OoDD performance than Claude 3.5 Sonnet and InternVL-76B. Claude 3.5 Sonnet frequently generates invalid responses compared to other proprietary models and struggles to detect near-OoD, resulting in worse performance on NINCO and SSB-Hard compared to the open-source model GLM-4v in terms of AUROC, despite achieving $16.58\%$ higher ID accuracy. InternVL2-26B achieves the best ID accuracy, but has difficulty with OoDD and shows the lowest valid response rate. While there is no clear-cut relationship between image recognition and OoDD capabilities, models with better image recognition performance generally exhibit stronger OoDD performance, consistent with the findings of Vaze et al. (2022).
@@
It is worth noting that the proprietary models generally perform on par with or better than the single-modal SOTA OoDD models. In addition, GPT-4o and Gemini Pro 1.5 outperform OpenCLIP in both ID classification and OoDD, showing significantly better performance in OoDD. For a more rigorous comparison, we evaluate OpenCLIP on the images where GPT-4o generates valid responses, as GPT-4o shows a relatively lower valid response rate. On the GPT-4o valid query set, GPT-4o still yields better results on both tasks. This demonstrates its superior visual recognition capability and its ability to express confidence scores through its generated responses. Detailed results can be found in Appendix B.2.
@@
Among the open-source models, one notable observation is their FPR of $100\%$. Our analysis indicates that this is caused by extremely biased confidence values, which will be further discussed in Sec. 3.6. GLM-4v exhibits the best OoDD performance, while InternVL2-76B shows the second-best zero-shot image recognition performance among the open-source models. Both models also demonstrate a better understanding of instructions than GPT-4o and Claude 3.5 Sonnet. In overall, InternVL2-76B shows superior performance in all aspects, including valid response rate, ID accuracy, and OoDD performance. It achieves higher ID accuracy and comparable OoDD performance to the single-modal SOTA models, alongside the highest valid response rates among all compared models. LLaVA-v1.6 struggles with image recognition, achieving an accuracy of just $2.25\%$. Most responses from LLaVA-v1.6 incorrectly identify ID images as OoD, resulting in poor classification accuracy and OoDD performance.

### v2

@@ Abstract
With the recent emergence of foundation models trained on internet-scale data and demonstrating remarkable generalization capabilities, such foundation models have become more widely adopted, leading to an expanding range of application domains. Despite this rapid proliferation, the trustworthiness of foundation models remains underexplored. Specifically, the out-of-distribution detection (OoDD) capabilities of large vision-language models (LVLMs), such as GPT-4o, which are trained on massive multi-modal data, have not been sufficiently addressed. The disparity between their demonstrated potential and practical reliability raises concerns regarding the safe and trustworthy deployment of foundation models. To address this gap, we evaluate and analyze the OoDD capabilities of various proprietary and open-source LVLMs. Our investigation contributes to a better understanding of how these foundation models represent confidence scores through their generated natural language responses. Furthermore, we propose a self-guided prompting approach, termed Reflexive Guidance (ReGuide), aimed at enhancing the OoDD capability of LVLMs by leveraging self-generated image-adaptive concept suggestions. Experimental results demonstrate that our ReGuide enhances the performance of current LVLMs in both image classification and OoDD tasks. [^1] https://github.com/daintlab/ReGuide
@@ 1 Introduction
Figure 1: Performance of proprietary and open-source LVLMs on the ImageNet200 benchmark: (a) state-of-the-art comparison and (b) the boosting effect of our ReGuide
Thanks to substantial advancements in hardware and the availability of large-scale datasets, foundation models have achieved remarkable performance across a wide range of tasks, demonstrating exceptional generalization. This evolution has shifted toward leveraging multiple data modalities, particularly vision and natural language. As a result, vision-language foundation models have demonstrated their capabilities across diverse domains, from general tasks like text-to-image generation and vision-question answering, to specialized fields such as medical diagnosis and robotics (Esser et al., 2024; Li et al., 2023; Wake et al., 2023; Majumdar et al., 2024). Despite the widespread adoption of these large vision-language models (LVLMs) including GPT (OpenAI, 2024), Claude (Anthropic, 2024), and Gemini (Reid et al., 2024), their trustworthiness and reliability have not been adequately investigated. Ensuring the robustness of deep neural networks has been a key research area to guarantee their safe application in practice. With the rapid popularization of LVLMs, practical concerns such as harmful content filtering and domain generalization have been actively studied (Zhang et al., 2024b; Han et al., 2024). However, fundamental aspects, such as the quality of confidence estimates for model predictions—extensively studied in single-modal models—remain underexplored. Some works have studied OoDD in CLIP (Ming et al., 2022; Jiang et al., 2024; Cao et al., 2024) and Stable Diffusion (Zhu et al., 2024), but few have examined OoDD in LVLMs, particularly with respect to confidence scores expressed through their natural language responses. This gap between their demonstrated capabilities and real-world reliability raises concerns about ensuring the safe and dependable deployment of LVLMs.
To bridge the gap, we first evaluate and compare the OoDD capabilities of LVLMs. The OoDD task allows us to investigate how LVLMs behave when required to generate responses beyond the categories defined within the user-provided prompt. Due to the lack of prior experimental configurations, we develop a framework for evaluating the OoDD capabilities of LVLMs. Our focus is on detecting OoD image inputs based on the in-distribution (ID) space specified by text inputs. To this end, we design a prompt that defines the ID space and guides the LVLM to provide confidence estimates through its generated responses. Using this prompt, we evaluate the OoDD capabilities of both proprietary and open-source LVLMs and analyze their behavior in expressing confidence from various perspectives. Fig. 1(a) summarizes the comparison of the evaluated LVLMs. In general, proprietary models outperform open-source models, and open-source models show performance improvements as model size increases. However, some open-source models including LLaVa-v1.6 (Mistral-7B) (Li et al., 2024a) and GLM-4v-9B (GLM et al., 2024) exhibit low OoDD performance despite achieving decent results on popular VLM benchmarks. [^2] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard We found that one of the reasons for this discrepancy is the insufficient image interpretation capabilities of these models, which limits their ability to accurately distinguish between ID and OoD inputs.
To enhance the OoD detectability of LVLMs, we propose a two-stage self-guided prompting approach called Reflexive Guidance (ReGuide). In the first stage, ReGuide prompts an LVLM to suggest two groups of concepts based on the given image: semantically similar concepts (i.e., near-OoD) and semantically dissimilar concepts (i.e., far-OoD). In the second stage, these suggested concepts are employed as auxiliary OoD classes. To the best of our knowledge, this is the first study to leverage image inputs to generate informative texts for OoDD. By utilizing the visual interpretation capabilities of LVLMs, ReGuide remains simple and model-agnostic, allowing the same prompt to be applied to different LVLMs. Fig. 1(b) shows the effectiveness of ReGuide. Notably, ReGuide significantly boosts the overall performance of open-source models, making them comparable to proprietary models. GPT-4o, which is the top performer prior to applying ReGuide, also benefits from ReGuide, especially on near-OoD datasets (e.g., NINCO, SSB Hard). Our results highlight the effectiveness of guiding LVLMs through self-generated, image-adaptive concept suggestions.
@@ 2 Related work
In single-modal vision models, OoDD has been actively studied. Starting with the baseline method of using the maximum softmax value as an OoD score (Hendrycks & Gimpel, 2017), methods have evolved to improve confidence modeling (Moon et al., 2020; Liu et al., 2020; Bibas et al., 2021), post-hoc techniques (Liang et al., 2018; Lee et al., 2018b; Sun et al., 2021; Djurisic et al., 2023; Sun et al., 2022; Zhang et al., 2023; Liu & Qin, 2024), or a combination of both (Xu et al., 2024). Another approach involves leveraging auxiliary OoD samples to better distinguish between ID and OoD inputs. While real OoD samples show strong detection performance (Hendrycks et al., 2019; Chen et al., 2021; Lu et al., 2023; Katz-Samuels et al., 2022), they require access to real OoD data. Synthetic samples offer an alternative, providing the benefits of auxiliary OoD data without the need for collecting real samples (Lee et al., 2018a; Du et al., 2022; Tao et al., 2023; Zheng et al., 2023).
@@
(a) Image classifier
(b) CLIP
(c) LVLM
Figure 2: Comparison of the OoDD framework for single-modal classifiers, CLIP, and LVLMs
@@ 3.1 Problem Definition
We frame the OoDD problem for LVLMs based on the zero-shot OoDD scenario defined for CLIP. Similarly, $\mathcal{Y}$ is set to a fixed word set containing only ID class words. Given both $\mathbf{x}$ and $\mathcal{Y}$ as inputs, an LVLM is instructed to produce prediction results in a structured format. Then, class predictions and confidence estimates for $\mathbf{x}$ are extracted from the natural language responses by the LVLM. If the ground-truth label of $\mathbf{x}$ is not in $\mathcal{Y}$, the LVLM should provide low confidence scores for all $y_{i}$ in $\mathcal{Y}$ in its answers. Fig. 2 summarizes the OoDD frameworks for single-modal image classifiers, CLIP, and LVLMs. Single-modal classifiers detect OoD images based solely on outputs from image inputs. In contrast, CLIP and LVLMs leverage both image and text inputs, with CLIP combining them via similarity measures and LVLMs integrating them through cross-modal interactions. LVLMs also generate natural language outputs, controllable through prompt instructions.
3.2 Prompt Design
Fig. 3 illustrates the framework for OoDD evaluation on LVLMs with a simplified prompt. Our prompt consists of four components: a task description, an explanation of the rejection class, guidelines, and examples for the response format. Unlike previous work on LVLMs (Han et al., 2024; Hwang et al., 2024; Groot & Valdenegro-Toro, 2024), we encounter a significant number of failure cases when using a simple prompt consisting of the task statement including $\mathcal{Y}$ and a formatted output structure. We attribute this inconsistency to differences in how confidence scores are assigned in our framework. In our prompt design, the LVLM is expected to provide confidence scores for each class in $\mathcal{Y}$, whereas prior work required only a single confidence score for the predicted class. To mitigate these failures, we enhance the prompt by adding the following components, resulting in the final prompt used in our experiments. The effectiveness of providing guidelines and examples and the complete prompt can be found in Appendix B.5 and B.6, respectively.
Rejection class. When we provide only $\mathcal{Y}$, we observe that predictions for OoD samples often do not correspond to any class in $\mathcal{Y}$. Due to its strong zero-shot visual recognition capabilities, the LVLM can either provide the ground-truth label for an OoD input image or reject classification into any of the classes in $\mathcal{Y}$. To address this unintended behavior, we introduce a rejection class named none of these classes and provide a detailed explanation of its role to the LVLM.
Guidelines addressing failure cases. In addition to the case where the prediction does not fall within $\mathcal{Y}$, we identify several other issues, such as mismatches between the class assigned the highest confidence score and the predicted class, assigning a confidence score of 0.0 to all classes including the rejection class, or failing to provide confidence scores for all classes in $\mathcal{Y}$. To reduce the frequency of these failures, we include guidelines that address the most common cases.
Response examples. In-context learning (ICL) is commonly used to improve response quality by providing models with examples that illustrate the desired format during inference, allowing them to generate outputs that better align with these examples (Brown et al., 2020). We leverage ICL to reduce failure caused by formatting issues. However, we observe that when only a single example is provided, the LVLM tends to mimic the example, regardless of whether the input image is ID or OoD. To mitigate this tendency, we provide examples for both ID and OoD input images.
@@ 3.4 Experimental Settings
Comparison models. To compare LVLMs from diverse perspectives, we consider both proprietary and open-source models. For proprietary models, we employ three state-of-the-art models: GPT-4o (2024-08-06) (OpenAI, 2024), Gemini Pro 1.5 (Reid et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024). For open-source models, we use four models: LLaVA-v1.6-Mistral-7B (LLaVA-v1.6) (Li et al., 2024a), GLM-4v-9B (GLM-4v) (GLM et al., 2024), QWEN-VL-Chat (QWEN) (Bai et al., 2023), InternVL2-InternLM2-Chat-26B (InternVL2-26B), and InternVL2-LLaMA3-76B (InternVL2-76B) (Chen et al., 2024). Additionally, we include OpenCLIP (ViT-B-32 pretrained on LAION 2B-s34b-b79k) (Cherti et al., 2023) as a non-generative LVLM. To further facilitate comparison between single- and multi-modal models, we also include three single-modal SOTA vision OoDD models, SCALE (Xu et al., 2024), fDBD (Liu & Qin, 2024), AugMix+ASH (Djurisic et al., 2023; Hendrycks et al., 2020).
@@
Evaluation metrics. We measure the OoDD performance using two commonly used metrics in OoDD: the area under the receiver operating characteristic curve (AUROC) and the false positive rate at the 95% true positive rate (FPR@95%TPR; FPR). We also include the ratio and number of valid responses—those that adhere to the provided guidelines and example format—as part of the evaluation metrics. We consider the ability to understand given instructions to be one of the model’s key capabilities. Only valid responses are included when measuring performance. We additionally evaluate models on the shared valid responses across all models for a rigorous comparison. This result can be found in Appendix B.2.
Table 1: Comparison on the ImageNet200 benchmark. Full model names are in the footnotes. ‘Valid’ indicates the ratio of valid responses out of a total of $23,031$ image-prompt queries, with counts in brackets. Bold highlights the best performance among generative LVLMs.
@@ 3.5 Results
Tab. 1 presents the OoDD capabilities of the compared models on the ImageNet200 benchmark. The near- and far-OoD results for SCALE and fDBD represent the average performance across their respective categories. ‘All OoD’ refers to the performance in distinguishing all OoD inputs, including near- and far-OoD inputs, from ID inputs (i.e., ID vs. all OoD). The results of QWEN-VL-Chat are ommitted due to its exceptionally low ability to follow instructions, with a valid response rate of less than $1\%$.
@@
On both image recognition and OoDD tasks, proprietary models outperform the open-source models in most cases, with reasonable valid response rates. All compared models have more difficulty in detecting near-OoD than far-OoD. The overall performance ranking of the evaluated LVLMs generally aligns with the OpenVLM leaderboard [^4] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard, except for Gemini Pro 1.5. In our results, Gemini Pro 1.5 shows better ID accuracy and OoDD performance than Claude 3.5 Sonnet and InternVL-76B. Claude 3.5 Sonnet frequently generates invalid responses compared to other proprietary models and struggles to detect near-OoD, resulting in worse performance on NINCO and SSB-Hard compared to the open-source model GLM-4v in terms of AUROC, despite achieving $16.65\%$ higher ID accuracy. InternVL2-26B achieves the best ID accuracy, but has difficulty with OoDD and shows the lowest valid response rate. While there is no clear-cut relationship between image recognition and OoDD capabilities, models with better image recognition performance generally exhibit stronger OoDD performance, consistent with the findings of Vaze et al. (2022).
@@
Notably, the proprietary models generally perform on par with or better than the single-modal SOTA OoDD models. In addition, GPT-4o and Gemini Pro 1.5 outperform OpenCLIP in both ID classification and OoDD, showing significantly better performance in OoDD. For a more rigorous comparison, we evaluate OpenCLIP on images where GPT-4o generates valid responses, as GPT-4o has a relatively lower valid response rate. On the GPT-4o valid query set, GPT-4o still yields better results on both tasks. This demonstrates its superior visual recognition capability and ability to express confidence scores through its generated responses. This observation is consistent with the evaluation on the shared valid set across all models compared. Detailed results can be found in Appendix B.2.
@@
Among the open-source models, one notable observation is their FPR of $100\%$. Our analysis indicates that this is caused by extremely biased confidence values, which will be further discussed in Sec. 3.6. GLM-4v exhibits the best OoDD performance, while InternVL2-76B shows the second-best zero-shot image recognition performance among the open-source models. Both models also demonstrate a better understanding of instructions than Claude 3.5 Sonnet. Overall, InternVL2-76B shows superior performance in all aspects, including valid response rate, ID accuracy, and OoDD performance. It achieves higher ID accuracy and comparable OoDD performance to the single-modal SOTA models, alongside the highest valid response rates among all compared models. LLaVA-v1.6 struggles with image recognition, achieving an accuracy of just $2.45\%$. Most responses from LLaVA-v1.6 incorrectly identify ID images as OoD, resulting in poor classification accuracy and OoDD performance.